OpenStudy (anonymous):

someone show work? Solve using substitution. y = 2x + 3 5x - y = -3

jhonyy9 (jhonyy9):

so i like help you but you need cooperat anything ,ok ?

jhonyy9 (jhonyy9):

how do you think what mean ,,substitution" method to solve it ?

jhonyy9 (jhonyy9):

so this mean that calcule the x ,for example from the first equation and substitute this result value of x in place of x in second equation and hence will can calcule easy the value of y ok ?

OpenStudy (anonymous):

i took 5x-1(2x+3)=-3 and then i did the work and for x i got 0 is that right so far?

jhonyy9 (jhonyy9):

so 5x-2x-3=-3 3x=0 x=0 yes is right

jhonyy9 (jhonyy9):

and now you need to calcule the y value

OpenStudy (anonymous):

so 5(0)-y=-3 0-y =-3 -0 -0 -y=-3 so the anwser is y=3

jhonyy9 (jhonyy9):

y=3 is right so do you understand it sure right ?

OpenStudy (anonymous):

yes now i do thank you for your help!!

jhonyy9 (jhonyy9):

yw was my pleasure bye